# Hurwitz moduli varieties parameterizing pointed covers of an algebraic curve with a fixed monodromy group

### v3

@@ 4. Universal families of pointed covers with fixed monodromy group
This induces a holomorphic map $X^{an}/G=Z\to(X/G)^{an}$ (cf. [27, Théorème 1.1]). The continuous map $|X^{an}/G|\to|(X/G)^{an}|$ is a homeomorphism (cf. Lemma 2.5 of [21]). The algebraic variety $X/G$ is normal, so $(X/G)^{an}$ is a normal complex space (cf. [27, Proposition 2.1]). Normality implies maximality (cf. [8, § 2.29]) therefore the holomorphic homeomorphism $X^{an}/G\to(X/G)^{an}$ is a biholomorphic map [8, § 2.29]. ∎
@@
Let $p:C\to Y$ be a $G$-cover. Let $b\in Y$ be a branch point. Let $p^{-1}(b)=\{w_{1},\ldots,w_{r}\}$. There is an embedded open disk $V\subset Y,b\in V$, such that $p^{-1}(V)=\sqcup_{i=1}^{r}W_{i}$ is a disjoint union of connected components $W_{i}$, $w_{i}\in W_{i}$ for $i=1,\ldots,r$, and every $p|_{W_{i}}:W_{i}\to V$ is a surjective cyclic analytic covering with Galois group $G(w_{i})=St_{G}(w_{i})\cong C_{e}\subset\mathbb{C}^{\ast}$, where $|G|=er$. Consider the left action of $G$ on $\Lambda\times C$ defined by $g(\lambda,z)=(\lambda g^{-1},gz)$. Let $X=(\Lambda\times C)/G:=\Lambda\times^{G}C$, let $\pi:\Lambda\times C\to X$ be the quotient map and let $f:X\to Y$ be the map $f(\pi(\lambda,z))=p(z)$. Let $i\in[1,r]$, $w=w_{i}$, $W=W_{i}$. It is clear that $f^{-1}(V)\cong(\Lambda\times W)/G(w)$. Let $x=\pi(\lambda,w)$. Let $\{\lambda_{1},\ldots,\lambda_{k}\}=\lambda G(w)$. Then ($(\{\lambda_{1},\ldots,\lambda_{k}\}\times W)/G(w)$ is the connected component of $f^{-1}(V)$ which contains $x=\pi(\lambda,w)$. Let $G(\lambda,w)=G(w)\cap G(\lambda)$. Let $|G(\lambda,w)|=q$. Then $e=kq$, the map ($(\{\lambda\}\times W)/G(\lambda,w)\to(\{\lambda_{1},\ldots,\lambda_{k}\}\times W)/G(w)$ is biholomorphic and the composition of holomorphic maps
@@
be the universal family of pointed $G$-covers of $(Y,y_{0})$ branched in $n$ points (cf. [21, Theorem 3.20]). We recall from [21, Section 3] the local analytic form of $p$ at the points lying over the branch locus $B=\{(y,(D,m))|y\in D\}$. Let $D=\{b_{1},\ldots,b_{j},\ldots,b_{n}\}$, $y_{0}\in Y\setminus D$. Let us choose local analytic coordinates $s_{i}$ at $b_{i}$, such that $s_{i}(b_{i})=0$, $i=1,\ldots,n$. Let $\epsilon\in\mathbb{R}^{+},\epsilon\ll 1$ be such that the open sets $U_{i}=\{y|s_{i}(y)<\epsilon\}$ have disjoint closures $\overline{U}_{i}$, $i=1,\ldots,n$ and $y_{0}\in Y\setminus\cup_{i=1}^{n}\overline{U}_{i}$. Let $\gamma_{1},\ldots,\gamma_{n}$ be closed paths based at $y_{0}$ as in § 2.6. Let $p(w)=(b,(D,m))$, where $b=b_{j}$. Let $U=U_{j}$ and let $V=U\times N_{(D,m)}(U_{1},\ldots,U_{n})$ (cf. § 3.4). For every $v=(y,(D^{\prime},m(D^{\prime}))\in V$, where $y\in U$, $D^{\prime}=\{y_{1},\ldots,y_{n}\},y_{i}\in U_{i}$ let $t_{i}(v)=s_{i}(y_{i})$ and let $t(v)=s_{j}(y)$. Let $G(w)$ be the isotropy group of $w$, $|G(w)|=e\geq 2$. Then $w$ has a connected neighborhood $W\subset|\mathcal{C}(y_{0})^{an}|$ which is $G(w)$-invariant, $p(W)=V$ and $gW\cap W=\emptyset$ if $g\in G\setminus G(w)$. Let $E\subset\mathbb{C}\times V$ be the analytic subset defined by $z^{e}=t-t_{j}$ and let $p_{1}:E\to V$ be the projection map. Then there is a biholomorphic map $\theta:W\to E$, such that $p_{1}\circ\theta=p|_{W}$. The composition $\rho=(z,t_{1},\ldots,t_{n})\circ\theta:W\to\mathbb{C}^{n+1}$ maps $W$ biholomorphically onto an open subset $\Omega\subset\mathbb{C}^{n+1}$. There exists a primitive character $\chi:G(w)\to\mathbb{C}^{\ast}$ such that $\theta$ and $\rho$ are $G(w)$-equivariant with respect to the action of $G(w)$ on $E$ and $\Omega$ defined respectively by $g(z,v)=(\chi(g)z,v)$ and $g(z,z_{1},\ldots,z_{n})=(\chi(g)z,z_{1},\ldots,z_{n})$.
@@
(13) $(\mathcal{O}_{\Omega,a})^{H(a)}=\psi^{\sharp}_{a}(\mathcal{O}_{\Omega_{1},b})$
Let $a=(a_{1},a_{2},\ldots,a_{n})$. If $a_{1}\neq 0$, then $H(a)=\{1\}$, $\psi:\Omega\to\Omega_{1}$ is locally biholomorphic at $a$, so (13) holds. Let $a_{1}=0$. Then $H(a)=H$. Without loss of generality we may suppose that $a=(0,0,\ldots,0)$. Let the $H$-invariant germ $f\in\mathcal{O}_{\Omega,a}$ be represented by a series $\sum_{\alpha_{1},\ldots,\alpha_{n}\geq 0}a_{\alpha_{1}\alpha_{2}\cdots\alpha_{n}}z_{1}^{\alpha_{1}}z_{2}^{\alpha_{2}}\cdots z_{n}^{\alpha_{n}}$ which converges absolutely in the polydisk $|z_{i}|<\varepsilon_{i}$, $i=1,\ldots,n$. One has $\alpha_{1}=q\beta_{1}$ for every $\alpha_{1}$. The series $\sum_{\beta_{1},\alpha_{2},\ldots,\alpha_{n}\geq 0}b_{\beta_{1}\alpha_{2}\cdots\alpha_{n}}y_{1}^{\beta_{1}}z_{2}^{\alpha_{2}}\cdots z_{n}^{\alpha_{n}}$ where $b_{\beta_{1}\alpha_{2}\cdots\alpha_{n}}=a_{q\beta_{1}\alpha_{2}\cdots\alpha_{n}}$ converges absolutely in the polydisk $|y_{1}|<\varepsilon_{1}^{q},|z_{i}|<\varepsilon_{i}$, $i=2,\ldots,n$ and represents a germ $f_{1}\in\mathcal{O}_{\Omega_{1},b}$ such that $f=\psi_{a}^{\sharp}(f_{1})$. Hence (13) holds for $\forall a\in\Omega$. ∎
In the next proposition we use the set-up and the notation of § 2.1 and § 4.4.
@@
It remains to prove the statements about the local analytic form of $\varphi$. According to Proposition 4.2 $\mathcal{X}(y_{0})^{an}$ is biholomorphic to $(\Lambda\times\mathcal{C}(y_{0})^{an})/G$. Let $G(\lambda,w)=G(\lambda)\cap G(w)$. This is a cyclic group of order $q$ and $e=kq$. The open subset $\{\lambda\}\times W\subset\Lambda\times\mathcal{C}(y_{0})$ is $G(\lambda,w)$-invariant and for every $g\in G\setminus G(\lambda,w)$ one has $g(\{\lambda\}\times W)\cap(\{\lambda\}\times W)=\emptyset$. Let $A=\pi(\{\lambda\}\times W)$. The open complex subspace $A\subset\mathcal{X}(y_{0})^{an}$ is biholomorphic to $\{\lambda\}\times W/G(\lambda,w)$. Let $H=G(\lambda,w)$ and let $\chi_{1}:H\to\mathbb{C}^{\ast}$ be the restriction of $\chi:G\to\mathbb{C}^{\ast}$ (cf. § 4.4). Identifying $\{\lambda\}\times W$ with $W$ the biholomorphic map $\theta:\{\lambda\}\times W\to E$ is $H$-equivariant. Let $\psi_{1}:E\to E_{1}$ be the map $\psi_{1}(z,v)=(z^{q},v)$. The composition $\psi_{1}\circ\theta$ is $H$-invariant, so there is a holomorphic map $\theta_{1}:A\to E_{1}$ such that $\psi_{1}\circ\theta=\theta_{1}\circ\pi|_{\{\lambda\}\times W}$. Let $\Omega=\rho(\{\lambda\}\times W)$, let $\psi:\mathbb{C}^{n+1}\to\mathbb{C}^{n+1}$ be the map $\psi(z,z_{1},\ldots,z_{n})=(z^{q},z_{1},\ldots,z_{n})$ and let $\Omega_{1}=\psi(\Omega)$. One has the following commutative diagram of holomorphic maps
@@
The vertical maps in the right square of (15) are $H$-invariant, the horizontal maps are biholomorphic (cf. [21, Proposition 3.18]) and by Lemma 4.5 $\Omega/H\cong\Omega_{1}$, therefore $E_{1}\cong E/H$. This implies that $\theta_{1}:A\to E_{1}$ and $\rho_{1}:A\to\Omega_{1}$ are biholomorphic maps. The map $p_{1}:E\to V$ equals $\varphi_{1}\circ\psi_{1}$, therefore the equality $p_{1}\circ\theta=p|_{\{\lambda\}\times W}$ implies $\varphi_{1}\circ\theta_{1}=\varphi_{A}$.
We mention that (ii) implies the smoothness of $\mathcal{X}(y_{0})$, as well as the smoothness of $\pi_{2}\circ\varphi:\mathcal{X}(y_{0})\to H^{G}_{n}(Y,y_{0})$ using [15, Ch. III Proposition 10.4]. ∎
@@
Let $(p:\mathcal{C}(y_{0})\to Y\times H^{G}_{n}(Y,y_{0}),\zeta)$ be as in § 4.4. Let $\rho:\mathcal{C}(y_{0})\to\mathcal{C}(y_{0})/G(\lambda_{0})$ be the quotient morphism and let $\eta_{1}=\rho\circ\zeta:H^{G}_{n}(Y,y_{0})\to\mathcal{C}(y_{0})/G(\lambda_{0})$. Then $(\overline{p}:\mathcal{C}(y_{0})/G(\lambda_{0})\to Y\times H^{G}_{n}(Y,y_{0}),\eta_{1})$ is a family of pointed covers of $(Y,y_{0})$ equivalent to the family $(\varphi:\mathcal{X}(y_{0})\to Y\times H^{G}_{n}(Y,y_{0}),\eta)$ defined in Proposition 4.6.
@@
The map $\delta:H^{G}_{n}(Y,y_{0})\to(Y\setminus y_{0})^{(n)}_{\ast}$ given by $\delta(D,m)=D$ is a topological covering map with respect to the canonical complex topologies, every neighborhood $N_{D}(U_{1},\ldots,U_{n})$ of $D\in(Y\setminus y_{0})^{(n)}_{\ast}$ is evenly covered (cf. (9)). It is clear from (8) that for every $\sigma\in N_{S(\Lambda)}(G)$ the map $(E,\mu)\mapsto(E,\sigma\mu\sigma^{-1})$ transforms $N_{(D,m)}(U_{1},\ldots,U_{n})$ into $N_{(D,\sigma m\sigma^{-1})}(U_{1},\ldots,U_{n})$, so $N_{S(\Lambda)}(G)$ acts on $H^{G}_{n}(Y,y_{0})$ by covering homeomorphism with respect to $|\delta^{an}|$. By [21, Corollary 4.5] this action is by covering automorphisms of the étale cover $\delta:H^{G}_{n}(Y,y_{0})\to(Y\setminus y_{0})^{(n)}_{\ast}$. The subgroup $N(\lambda_{0})$ acts freely on $H^{G}_{n}(Y,y_{0})$ (cf. Proposition 2.8), so the quotient set $H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})=H^{G}_{n}(Y,y_{0})/N(\lambda_{0})$ has a structure of a smooth, affine algebraic variety [28, Ch.III Prop.18] and the quotient map $\nu$ is an étale cover. One has $\delta=\delta_{1}\circ\nu$, so $\delta_{1}$ is a morphism by the universal property of quotients and it is an étale cover since $\delta$ has this property and $\nu$ is étale. ∎
@@
The quotient morphism $\nu:H^{G}_{n}(Y,y_{0})\to H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$ yields a homeomorphism $|H^{G}_{n}(Y,y_{0})^{an}|/N(\lambda_{0})\overset{\sim}{\longrightarrow}|H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})^{an}|$ (cf. [21, Lemma 2.5]), so $|H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})^{an}|$ has a neighborhood basis consisting of the open sets
@@
Let $\mathcal{C}(y_{0})^{\prime}=p^{-1}(Y\times H^{G}_{n}(Y,y_{0})\setminus B)$. We recall from [21, Section 3] that $\mathcal{C}(y_{0})^{\prime}$ is bijective to the set $\{(\Gamma_{m}[\alpha]_{D},D,m)\}$ where $(D,m)\in H^{G}_{n}(Y,y_{0})$, $\Gamma_{m}=\operatorname{Ker}(m:\pi_{1}(Y\setminus D,y_{0})\to G)$, $\alpha$ is a path in $Y\setminus D$ with $\alpha(0)=y_{0}$ and $[\alpha]_{D}$ is its homotopy class in $Y\setminus D$. The map $p^{\prime}=p|_{\mathcal{C}(y_{0})^{\prime}}:\mathcal{C}(y_{0})^{\prime}\to Y\times H^{G}_{n}(Y,y_{0})\setminus B$, defined by $(\Gamma_{m}[\alpha]_{D},D,m)\mapsto(\alpha(1),(D,m))$, is a topological Galois covering map with respect to the topologies of the associated complex spaces, with group of Deck transformations isomorphic to $G$, where the action of $G$ is defined as follows: if $g\in G$, $g=m([\sigma]_{D})$, then $g(\Gamma_{m}[\alpha]_{D},D,m)=(\Gamma_{m}[\sigma\cdot\alpha]_{D},D,m)$.
In the set-up of § 2.1 for every $(D,m)\in H^{G}_{n}(Y,y_{0})$ let $\Gamma_{m,\lambda_{0}}=m^{-1}(G(\lambda_{0}))\subset\pi_{1}(Y\setminus D,y_{0})$. Let $\mathcal{X}(y_{0})^{\prime}=\varphi^{-1}(Y\times H^{G}_{n}(Y,y_{0})\setminus B)$. By Proposition 4.7 $\mathcal{X}(y_{0})^{\prime}$ is isomorphic to $\mathcal{C}(y_{0})^{\prime}/G(\lambda_{0})$ by a covering isomorphism over $Y\times H^{G}_{n}(Y,y_{0})\setminus B$. The quotient $\mathcal{C}(y_{0})^{\prime}/G(\lambda_{0})$ is bijective to the set $\{(\Gamma_{m,\lambda_{0}}[\alpha]_{D},D,m)\}$ and the quotient morphism $\rho^{\prime}:\mathcal{C}(y_{0})^{\prime}\to\mathcal{C}(y_{0})^{\prime}/G(\lambda_{0})$ is given by $(\Gamma_{m}[\alpha]_{D},D,m)\mapsto(\Gamma_{m,\lambda_{0}}[\alpha]_{D},D,m)$. One has ${\mathcal{X}(y_{0})^{\prime}}^{an}\cong{\mathcal{C}(y_{0})^{\prime}}^{an}/G(\lambda_{0})$ by Proposition 4.2. The map $\varphi^{\prime}:\mathcal{X}(y_{0})^{\prime}\to Y\times H^{G}_{n}(Y,y_{0})\setminus B$ may be identified with
@@
[…]

### v4

@@ 4. Universal families of pointed covers with fixed monodromy group
This induces a holomorphic map $X^{an}/G=Z\to(X/G)^{an}$ (cf. [28, Théorème 1.1]). The continuous map $|X^{an}/G|\to|(X/G)^{an}|$ is a homeomorphism (cf. Lemma 2.5 of [22]). The algebraic variety $X/G$ is normal, so $(X/G)^{an}$ is a normal complex space (cf. [28, Proposition 2.1]). Normality implies maximality (cf. [8, §2.29]) therefore the holomorphic homeomorphism $X^{an}/G\to(X/G)^{an}$ is a biholomorphic map [8, §2.29]. ∎
@@
Let $p:C\to Y$ be a $G$-cover. Let $b\in Y$ be a branch point. Let $p^{-1}(b)=\{w_{1},\ldots,w_{r}\}$. There is an embedded open disk $V\subset Y,b\in V$, such that $p^{-1}(V)=\sqcup_{i=1}^{r}W_{i}$ is a disjoint union of connected components $W_{i}$, $w_{i}\in W_{i}$ for $i=1,\ldots,r$, and every $p|_{W_{i}}:W_{i}\to V$ is a surjective cyclic analytic covering with Galois group $G(w_{i})=St_{G}(w_{i})\cong C_{e}\subset\mathbb{C}^{\ast}$, where $|G|=er$. Consider the left action of $G$ on $\Lambda\times C$ defined by $g(\lambda,z)=(\lambda g^{-1},gz)$. Let $X=(\Lambda\times C)/G:=\Lambda\times^{G}C$, let $\pi:\Lambda\times C\to X$ be the quotient map and let $f:X\to Y$ be the map $f(\pi(\lambda,z))=p(z)$. Let $i$ be an integer, $1\leq i\leq r$, let $w=w_{i}$, $W=W_{i}$. It is clear that $f^{-1}(V)\cong(\Lambda\times W)/G(w)$. Let $x=\pi(\lambda,w)$. Let $\{\lambda_{1},\ldots,\lambda_{k}\}=\lambda G(w)$. Then $(\{\lambda_{1},\ldots,\lambda_{k}\}\times W)/G(w)$ is the connected component of $f^{-1}(V)$ which contains $x=\pi(\lambda,w)$. Let $G(\lambda,w)=G(w)\cap G(\lambda)$. Let $|G(\lambda,w)|=q$. Then $e=kq$, the map $(\{\lambda\}\times W)/G(\lambda,w)\to(\{\lambda_{1},\ldots,\lambda_{k}\}\times W)/G(w)$ is biholomorphic and the composition of holomorphic maps
@@
be the universal family of pointed $G$-covers of $(Y,y_{0})$ branched in $n$ points (cf. [22, Theorem 3.20]). We recall from [22, Section 3] the local analytic form of $p$ at the points lying over the branch locus $B=\{(y,(D,m))|y\in D\}$. Let $D=\{b_{1},\ldots,b_{j},\ldots,b_{n}\}$, $y_{0}\in Y\setminus D$. Let us choose local analytic coordinates $s_{i}$ at $b_{i}$, such that $s_{i}(b_{i})=0$, $i=1,\ldots,n$. Let $\epsilon\in\mathbb{R}^{+},\epsilon\ll 1$ be such that the open sets $U_{i}=\{y|s_{i}(y)<\epsilon\}$ have disjoint closures $\overline{U}_{i}$, $i=1,\ldots,n$ and $y_{0}\in Y\setminus\cup_{i=1}^{n}\overline{U}_{i}$. Let $\gamma_{1},\ldots,\gamma_{n}$ be closed paths based at $y_{0}$ as in §2.6. Let $p(w)=(b,(D,m))$, where $b=b_{j}$. Let $U=U_{j}$ and let $V=U\times N_{(D,m)}(U_{1},\ldots,U_{n})$ (cf. §3.4). For every $v=(y,(D^{\prime},m(D^{\prime}))\in V$, where $y\in U$, $D^{\prime}=\{y_{1},\ldots,y_{n}\},y_{i}\in U_{i}$ let $t_{i}(v)=s_{i}(y_{i})$ and let $t(v)=s_{j}(y)$. Let $G(w)$ be the isotropy group of $w$, $|G(w)|=e\geq 2$. Then $w$ has a connected neighborhood $W\subset|\mathcal{C}(y_{0})^{an}|$ which is $G(w)$-invariant, $p(W)=V$ and $gW\cap W=\emptyset$ if $g\in G\setminus G(w)$. Let $E\subset\mathbb{C}\times V$ be the analytic subset defined by $z^{e}=t-t_{j}$ and let $p_{1}:E\to V$ be the projection map. Then there is a biholomorphic map $\theta:W\to E$, such that $p_{1}\circ\theta=p|_{W}$. The composition $\rho=(z,t_{1},\ldots,t_{n})\circ\theta:W\to\mathbb{C}^{n+1}$ maps $W$ biholomorphically onto an open subset $\Omega\subset\mathbb{C}^{n+1}$. There exists a primitive character $\chi:G(w)\to\mathbb{C}^{\ast}$ such that $\theta$ and $\rho$ are $G(w)$-equivariant with respect to the action of $G(w)$ on $E$ and $\Omega$ defined respectively by $g(z,v)=(\chi(g)z,v)$ and $g(z,z_{1},\ldots,z_{n})=(\chi(g)z,z_{1},\ldots,z_{n})$.
@@
(13) $(\mathcal{O}_{\Omega,a})^{H(a)}=\psi^{\sharp}_{a}(\mathcal{O}_{\Omega_{1},b}).$
Let $a=(a_{1},a_{2},\ldots,a_{n})$. If $a_{1}\neq 0$, then $H(a)=\{1\}$, $\psi:\Omega\to\Omega_{1}$ is locally biholomorphic at $a$, so (13) holds. Let $a_{1}=0$. Then $H(a)=H$. Without loss of generality we may suppose that $a=(0,0,\ldots,0)$. Let the $H$-invariant germ $f\in\mathcal{O}_{\Omega,a}$ be represented by a series $\sum_{\alpha_{1},\ldots,\alpha_{n}\geq 0}a_{\alpha_{1}\alpha_{2}\cdots\alpha_{n}}z_{1}^{\alpha_{1}}z_{2}^{\alpha_{2}}\cdots z_{n}^{\alpha_{n}}$ which converges absolutely in the polydisk $|z_{i}|<\varepsilon_{i}$, $i=1,\ldots,n$. One has $\alpha_{1}=q\beta_{1}$ for every $\alpha_{1}$. The series $\sum_{\beta_{1},\alpha_{2},\ldots,\alpha_{n}\geq 0}b_{\beta_{1}\alpha_{2}\cdots\alpha_{n}}y_{1}^{\beta_{1}}z_{2}^{\alpha_{2}}\cdots z_{n}^{\alpha_{n}}$, where $b_{\beta_{1}\alpha_{2}\cdots\alpha_{n}}=a_{q\beta_{1}\alpha_{2}\cdots\alpha_{n}}$, converges absolutely in the polydisk $|y_{1}|<\varepsilon_{1}^{q},|z_{i}|<\varepsilon_{i}$, $i=2,\ldots,n$ and represents a germ $f_{1}\in\mathcal{O}_{\Omega_{1},b}$ such that $f=\psi_{a}^{\sharp}(f_{1})$. Hence (13) holds for $\forall a\in\Omega$. ∎
In the next proposition we use the setup and the notation of §2.1 and §$4.$
@@
It remains to prove the statements about the local analytic form of $\varphi$. According to Proposition 4.2 $\mathcal{X}(y_{0})^{an}$ is biholomorphic to $(\Lambda\times\mathcal{C}(y_{0})^{an})/G$. Let $G(\lambda,w)=G(\lambda)\cap G(w)$. This is a cyclic group of order $q$ and $e=kq$. The open subset $\{\lambda\}\times W\subset\Lambda\times\mathcal{C}(y_{0})$ is $G(\lambda,w)$-invariant and for every $g\in G\setminus G(\lambda,w)$ one has $g(\{\lambda\}\times W)\cap(\{\lambda\}\times W)=\emptyset$. Let $A=\pi(\{\lambda\}\times W)$. The open complex subspace $A\subset\mathcal{X}(y_{0})^{an}$ is biholomorphic to $\{\lambda\}\times W/G(\lambda,w)$. Let $H=G(\lambda,w)$ and let $\chi_{1}:H\to\mathbb{C}^{\ast}$ be the restriction of $\chi:G\to\mathbb{C}^{\ast}$ (cf. §4.4). Identifying $\{\lambda\}\times W$ with $W$ the biholomorphic map $\theta:\{\lambda\}\times W\to E$ is $H$-equivariant. Let $\psi_{1}:E\to E_{1}$ be the map $\psi_{1}(z,v)=(z^{q},v)$. The composition $\psi_{1}\circ\theta$ is $H$-invariant, so there is a holomorphic map $\theta_{1}:A\to E_{1}$ such that $\psi_{1}\circ\theta=\theta_{1}\circ\pi|_{\{\lambda\}\times W}$. Let $\Omega=\rho(\{\lambda\}\times W)$, let $\psi:\mathbb{C}^{n+1}\to\mathbb{C}^{n+1}$ be the map $\psi(z,z_{1},\ldots,z_{n})=(z^{q},z_{1},\ldots,z_{n})$ and let $\Omega_{1}=\psi(\Omega)$. One has the following commutative diagram of holomorphic maps:
@@
The vertical maps in the right square of (15) are $H$-invariant, the horizontal maps are biholomorphic (cf. [22, Proposition 3.18]) and by Lemma 4.5 $\Omega/H\cong\Omega_{1}$, therefore $E_{1}\cong E/H$. This implies that $\theta_{1}:A\to E_{1}$ and $\rho_{1}:A\to\Omega_{1}$ are biholomorphic maps. The map $p_{1}:E\to V$ equals $\varphi_{1}\circ\psi_{1}$, therefore the equality $p_{1}\circ\theta=p|_{\{\lambda\}\times W}$ implies $\varphi_{1}\circ\theta_{1}=\varphi|_{A}$.
We mention that (ii) implies the smoothness of $\mathcal{X}(y_{0})$, as well as the smoothness of $\pi_{2}\circ\varphi:\mathcal{X}(y_{0})\to H^{G}_{n}(Y,y_{0})$ using [16, Ch. III Proposition 10.4]. ∎
@@
Let $(p:\mathcal{C}(y_{0})\to Y\times H^{G}_{n}(Y,y_{0}),\zeta)$ be as in §$4.$ Let $\rho:\mathcal{C}(y_{0})\to\mathcal{C}(y_{0})/G(\lambda_{0})$ be the quotient morphism and let $\eta_{1}=\rho\circ\zeta:H^{G}_{n}(Y,y_{0})\to\mathcal{C}(y_{0})/G(\lambda_{0})$. Then $(\overline{p}:\mathcal{C}(y_{0})/G(\lambda_{0})\to Y\times H^{G}_{n}(Y,y_{0}),\eta_{1})$ is a family of pointed covers of $(Y,y_{0})$ equivalent to the family $(\varphi:\mathcal{X}(y_{0})\to Y\times H^{G}_{n}(Y,y_{0}),\eta)$ defined in Proposition 4.6.
@@
The map $\delta:H^{G}_{n}(Y,y_{0})\to(Y\setminus y_{0})^{(n)}_{\ast}$ given by $\delta(D,m)=D$ is a topological covering map with respect to the canonical complex topologies, every neighborhood $N_{D}(U_{1},\ldots,U_{n})$ of $D\in(Y\setminus y_{0})^{(n)}_{\ast}$ is evenly covered (cf. (9)). It is clear from (8) that for every $\sigma\in N_{S(\Lambda)}(G)$ the map $(E,\mu)\mapsto(E,\sigma\mu\sigma^{-1})$ transforms $N_{(D,m)}(U_{1},\ldots,U_{n})$ into $N_{(D,\sigma m\sigma^{-1})}(U_{1},\ldots,U_{n})$, so $N_{S(\Lambda)}(G)$ acts on $H^{G}_{n}(Y,y_{0})$ by covering homeomorphism with respect to $|\delta^{an}|$. By [22, Corollary 4.5] this action is by covering automorphisms of the étale cover $\delta:H^{G}_{n}(Y,y_{0})\to(Y\setminus y_{0})^{(n)}_{\ast}$. The subgroup $N(\lambda_{0})$ acts freely on $H^{G}_{n}(Y,y_{0})$ (cf. Proposition 2.8), so the quotient set $H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})=H^{G}_{n}(Y,y_{0})/N(\lambda_{0})$ has a structure of a smooth, affine algebraic variety [29, Ch.III Prop.18] and the quotient map $\nu$ is an étale cover. One has $\delta=\delta_{1}\circ\nu$, so $\delta_{1}$ is a morphism by the universal property of quotients and it is an étale cover since $\delta$ has this property and $\nu$ is étale. ∎
@@
The quotient morphism $\nu:H^{G}_{n}(Y,y_{0})\to H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})$ yields a homeomorphism $|H^{G}_{n}(Y,y_{0})^{an}|/N(\lambda_{0})\overset{\sim}{\longrightarrow}|H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})^{an}|$ (cf. [22, Lemma 2.5]), so $|H^{\Lambda,G}_{n,\lambda_{0}}(Y,y_{0})^{an}|$ has a neighborhood basis consisting of the open sets
@@
Let $\mathcal{C}(y_{0})^{\prime}=p^{-1}(Y\times H^{G}_{n}(Y,y_{0})\setminus B)$. We recall from [22, Section 3] that $\mathcal{C}(y_{0})^{\prime}$ is bijective to the set $\{(\Gamma_{m}[\alpha]_{D},D,m)\}$ where $(D,m)\in H^{G}_{n}(Y,y_{0})$, $\Gamma_{m}=\operatorname{Ker}(m:\pi_{1}(Y\setminus D,y_{0})\to G)$, $\alpha$ is a path in $Y\setminus D$ with $\alpha(0)=y_{0}$ and $\alpha]_{D}$ is its homotopy class in $Y\setminus D$. The map $p^{\prime}=p|_{\mathcal{C}(y_{0})^{\prime}}:\mathcal{C}(y_{0})^{\prime}\to Y\times H^{G}_{n}(Y,y_{0})\setminus B$, defined by $(\Gamma_{m}[\alpha]_{D},D,m)\mapsto(\alpha(1),(D,m))$, is a topological Galois covering map with respect to the topologies of the associated complex spaces, with group of Deck transformations isomorphic to $G$, where the action of $G$ is defined as follows: if $g\in G$, $g=m([\sigma]_{D})$, then $g(\Gamma_{m}[\alpha]_{D},D,m)=(\Gamma_{m}[\sigma\cdot\alpha]_{D},D,m)$.
In the setup of §2.1 for every $(D,m)\in H^{G}_{n}(Y,y_{0})$ let $\Gamma_{m,\lambda_{0}}=m^{-1}(G(\lambda_{0}))\subset\pi_{1}(Y\setminus D,y_{0})$. Let $\mathcal{X}(y_{0})^{\prime}=\varphi^{-1}(Y\times H^{G}_{n}(Y,y_{0})\setminus B)$. By Proposition 4.7 $\mathcal{X}(y_{0})^{\prime}$ is isomorphic to $\mathcal{C}(y_{0})^{\prime}/G(\lambda_{0})$ by a covering isomorphism over $Y\times H^{G}_{n}(Y,y_{0})\setminus B$. The quotient $\mathcal{C}(y_{0})^{\prime}/G(\lambda_{0})$ is bijective to the set $\{(\Gamma_{m,\lambda_{0}}[\alpha]_{D},D,m)\}$ and the quotient morphism $\rho^{\prime}:\mathcal{C}(y_{0})^{\prime}\to\mathcal{C}(y_{0})^{\prime}/G(\lambda_{0})$ is given by $(\Gamma_{m}[\alpha]_{D},D,m)\mapsto(\Gamma_{m,\lambda_{0}}[\alpha]_{D},D,m)$. One has ${\mathcal{X}(y_{0})^{\prime}}^{an}\cong{\mathcal{C}(y_{0})^{\prime}}^{an}/G(\lambda_{0})$ by Proposition 4.2. The map $\varphi^{\prime}:\mathcal{X}(y_{0})^{\prime}\to Y\times H^{G}_{n}(Y,y_{0})\setminus B$ may be identified with
@@
$\mathcal{X}(y_{0})\overset{\varphi}{\longrightarrow}Y\times H^{G}_{n}(Y,y_{0})\overset{id\times\delta}{\longrightarrow}Y\times(Y\setminus y_{0})^{(n)}_{\ast},$